3x-4y=-9
-x-4y=-13

Answer by stanbon(57219)

3x-4y=-9
-x-4y=-13
----------------
Subtract the 2nd equation from the 1st to solve for "x":
4x = 4
x = 1
-------------------
Substitute into 3x-4y=-9 to solve for "y":
3*1 - 4y = -9
4y = 12
y = 3
-------------------
Cheers,
Stan H.
Reply to stanbon@comcast.net or algebra.com